OpenStudy (anonymous):

Could someone solve y=2x^2/x+1 at x=2 using the chain rule. I keep getting 8/9 as an answer, but its wrong and I don't know what I'm doing wrong.

OpenStudy (kropot72):

Is this the equation? \[y=\frac{2x ^{2}}{x+1}\]

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

Do you want dy/dx at x = 2 ?

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

\[if\ y=\frac{u}{v}\ then\ \frac{dy}{dx}=\frac{v\frac{du}{dx}-u\frac{dv}{dx}}{v ^{2}}\] In your case \[u=2x ^{2}\] \[v=x+1\] Can you carry on with the differentiation?

OpenStudy (anonymous):

WAIT!!!!! i know what I was doing wrong. Thank you

OpenStudy (kropot72):

Gr8! If you need more help please ask :)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

I need help with another question